# Long-term prediction of El Niño-Southern Oscillation using reservoir computing with data-driven realtime filter

### v1

@@ IV.1 Discussions
In complex phenomena, such as atmosphere-ocean coupled dynamics, multiple processes operating at various speeds are intertwined, forming cross-scale interactions. In such cases it is known that the prediction accuracy of the reservoir computing model varies depending on the passband of the filterHassanibesheli, Kurths, and Boers (2022); Huang, Yang, and Fu (2020). Figure 7 shows the power spectrum of the time series before and after applying the filter. It can be observed that the filter significantly reduce fluctuations with periods shorter than approximately three years and those longer than approximately eight years. For a practical methodology for prediction, it is necessary to identify filter parameter settings that maximize the model’s predictive performance while minimizing the deviation of the filtered time series from its original.
@@
Figure 7: Temporal characteristics of the original and filtered realtime SST anomaly. (a) Power spectrum and (b) autocorrelation. x axis of (a) indicates the frequency in cycles per month
@@
In this paper, fluctuations with periods shorter than one to two years are removed. It is suggested that the reservoir computing model can more easily capture the primary timescale variations of ENSO through this filtering process. By applying the filter, the local minimum at about 22 months and the local maximum at about 44 months of the autocorrelation of the realtime SST anomaly are amplified (Figure 7b). This implies that the periodic fluctuations of about 44 months are relatively strengthened, which in turn enhances the model’s predictive performance on the lead time of half of the period, namely, approximately 22 months.
@@
One characteristic observed in the all-season correlation skill of the reservoir model’s predictions is that, it decreases from unity until reaching a local minimum at around 15 months, and then increases again up to a lead time of 23 months (Figure 5). Similar temporary increase in correlation or decrease in relative error have been reported in previous studies using band-pass filtersHassanibesheli, Kurths, and Boers (2022); Huang and Fu (2023). More detailed investigation into the relationship between the prediction skills and the passband settings of the filter is required in our future study.
@@ Acknowledgements.
TJ was supported by University of Toyama Data Science Promotion Project. Research of TY was partly supported by the JSPS Grants-in-Aid for Scientific Research 24H00186. TM was supported by the JST Moonshot R&D (JPMJMS2389). KN was supported by JSPS KAKENHI Grant No.22K17965. YS was supported by JSPS KAKENHI Grant No.19KK0067, 21K18584, 23H04465 and 24K00537.

### v2

@@ IV.1 Discussions
In this study, we used a delay-coordinate vector of filtered realtime SST anomaly as input and output of the ESN model. According to the model construction guidelines provided by Nakai and Saiki (2021)Nakai and Saiki (2021), setting the value of the delay $\Delta\tau$ such that the consecutive two components of a delay-coordinate vector have a certain degree of correlation allows for capturing the fast dynamics. In addition, $\Delta\tau M$ should be chosen so that the envelope of the autocorrelation of the target time series is between 0.35 and 0.4 to capture the slow macroscopic dynamics. The envelope estimated from the filtered realtime SST anomaly used in this study (dotted line in Figure 7b) passes through this range between lags of 35 to 39 months. Therefore, the hyperparameter combination obtained through the Bayesian optimization (Table 2, $\Delta\tau M=36$) meets this criterion. By incorporating past information from up to 36 months before the forecast start time, it is inferred that the long-term dynamical information of ENSO is effectively captured.
In our estimation, the dimension of the reservoir state vector is $N=244$, and for several other good combinations of hyperparameters obtained through Bayesian optimization, $N$ typically ranges from 100 to 300. One limitation of our current approach is that, considering the long-term dynamics of ENSO of interest, the available length of time series data based on historical observations as training data is limited. In the current work, setting a large number of nodes in the reservoir does not offer a significant advantage in reproducing long-term dynamics. If longer data are obtained by some data augmentation, increasing the number of nodes may allow for constructing models that can express more complex dynamics.
@@
In complex phenomena, such as atmosphere-ocean coupled dynamics, multiple processes operating at various speeds are intertwined, forming cross-scale interactions. In such cases, it is known that the prediction accuracy of the reservoir computing model varies depending on the passband of the filterHassanibesheli, Kurths, and Boers (2022); Huang, Yang, and Fu (2020). Figure 7 shows the power spectrum of the time series before and after applying the filter. It can be observed that the filter significantly reduces fluctuations with periods shorter than approximately three years and those longer than approximately eight years. For a practical methodology for prediction, it is necessary to identify filter parameter settings that maximize the model’s predictive performance while minimizing the deviation of the filtered time series from its original.
@@
Figure 7: Temporal characteristics of the original and filtered realtime SST anomaly. (a) Power spectrum of the original (black line) and filtered (red line) time series. (b) Autocorrelation of the time series. Black dotted lines denote visually estimated envelope of filtered time series (±$\pm\exp[-t/37]$).
In this paper, fluctuations with periods shorter than one to two years are removed. Here, the same metric as in previous research is used for prediction skill, but the target time series is the filtered realtime SST anomaly, which is not identical to the conventional Niño 3.4 index. Therefore, a direct comparison of the prediction horizon to those of previous studies is not possible. However, since the correlation between the original and filtered time series is high, it can be said that similar time series are being predicted. When trying to accurately predict the most unstable high-frequency dynamics, the accuracy of long-term predictions tends to decrease. Thus, predicting the original realtime SST anomaly is not the aim of this study. However, considering that the all-season correlation skill gradually decreases from 1 in Fig. 5 and the high correlation between the original and filtered time series, it can be inferred that the reliability of short-term predictions for several months is not significantly compromised.
@@
Furthermore, from the prediction skill, it is possible to gain some insights into the time scales of ENSO dynamics that each model is particularly proficient in forecasting. In the model of Hassanibesheli et al. (2022)Hassanibesheli, Kurths, and Boers (2022), the correlation remains at a high value of around 0.8 until 14 months of lead time. The advantage of combining low-frequency variation prediction using an ESN with high-frequency variation prediction using a statistical method called “past noise forecasting” is reflected in the prediction skill at a timescale of approximately one year. In this study, it is suggested that the reservoir computing model can more easily capture the primary timescale variations of ENSO through this filtering process. By applying the filter, the local minimum at about 22 months and the local maximum at about 44 months of the autocorrelation of the realtime SST anomaly are amplified (Figure 7b). This implies that the periodic fluctuations of about 44 months are relatively enhanced, which in turn improves the model’s predictive performance on the lead time of half of the period, namely, approximately 22 months.
@@
One characteristic observed in the all-season correlation skill of the reservoir model’s predictions is that, it decreases from unity until reaching a local minimum at around 15 months, and then increases again up to a lead time of 23 months (Figure 5). Similar temporary increase in correlation or decrease in relative error have been reported in previous studiesHassanibesheli, Kurths, and Boers (2022); Huang and Fu (2023). The non-monotonic behavior in prediction skill is presumably due to the calculation method of the metric. As shown in Eq. (7), on calculating $C(\mu)$, only the value at the prediction lead time $\mu$ is used, rather than using the time series from lead time 1 to $\mu$. Therefore, it is possible for longer-term predictions to have higher scores than shorter-term predictions. A more detailed investigation into the relationship between the prediction skills and the passband settings of the filter is required in our future study.
Determining the optimal combination of hyperparameters for machine learning using a simple grid search becomes increasingly difficult as the number of hyperparameters grows. Considering the simplicity of the architecture and the large number of hyperparameters in reservoir computing, Bayesian optimization for tuning is an effective solution for improving its prediction performanceJoy, Mattheakis, and Protopapas (2022); Akiba et al. (2019). In this study, six parameters were tuned for the optimization of the filter, and nine parameters were tuned for the optimization of the reservoir model. In both cases, up to 1000 combination trials were conducted. The final performance did not change after about 100 trials for the filter, and after about 200 trials for the reservoir model. Therefore, we believe that the high dimensionality of the search space is not a crucial barrier for the specific task in this study. With the advancement of state-of-the-art optimization algorithms, it is expected that future research will be able to utilize more sophisticated reservoir models with an increased number of hyperparameters.
@@ Acknowledgements.
Research of TY was partly supported by the JSPS Grants-in-Aid for Scientific Research 24H00186. TM was supported by the JST Moonshot R&D (JPMJMS2389). KN was supported by JSPS KAKENHI Grant No.22K17965. YS was supported by JSPS KAKENHI Grant No.19KK0067, 21K18584, 23H04465 and 24K00537.